OpenStudy (anonymous):

If p + 3r is equal to 150 percent of 6r, what is the value of p/r ?

OpenStudy (anonymous):

p + 3r = 9r

OpenStudy (lgbasallote):

p + 3r is 150% of 6r so... \[ p + 3r = \frac{150}{100} \times 6r\] so... \[100(p + 3r) = 150 \times 6r\] \[\implies 100p + 300r = 900r\] \[\implies 100p = 900 r - 300r\] \[\implies 100p = 600r\] divide both sides by 600r \[\implies \frac{100p}{600r} = 1\] simplify \[\implies \frac {p}{6r} = 1\] cross multiply.. \[\implies \frac pr = 6\] does that help?

OpenStudy (anonymous):

sorry @lgbasallote site is weird

OpenStudy (lgbasallote):

yeah. so they say...

OpenStudy (anonymous):

Thankyou so much! it actually did!! Wow gee thanks its much appriciated!

OpenStudy (anonymous):

Are you sure though? Like completley positive??

OpenStudy (lgbasallote):

very positive. like pregnancy positive! :D

OpenStudy (anonymous):

LOL thanksss :)

OpenStudy (lgbasallote):

welcome ^_^